# Non-perturbative renormalization group for Higgs-like models in 4D

### v2

@@ Abstract
We recently defined a model of 2 coupled SU(2) Higgs-like doublets which revealed an interesting structure of renormalization group flows to 1-loop when the SU(2) is broken to U(1). In this article we compute the beta functions to 3 loops and show that the 1-loop structure of flows persists to higher orders. For SU(2) broken to U(1), we conjecture a beta function to all orders. The flows can be extended to large coupling using a strong-weak coupling duality $g\to 1/g$. One finds a line of fixed points which are new conformal field theories in 4 spacetime dimensions which are non-unitary due to negative norm states but still have real eigenvalues. We also find massless flows between 2 non-trivial fixed points, and a regime with a cyclic RG flow. For the flows between points on the critical line, we compute the anomalous dimensions of the perturbations in the UV and IR, and identify some special points where anomalous dimensions are rational numbers. The model is non-unitary since the hamiltonian is pseudo-hermitian, $H^{\dagger}={\cal K}H{\cal K}^{\dagger}$. The unitary operator ${\cal K}$ satisfies ${\cal K}^{2}=1$ and this allows a projection onto positive norm states with a unitary time evolution with positive probabilities.
@@ I Introduction
There is a conspicuous shortage of well-understood conformal field theories (CFT’s) in 4 spacetime dimensions that aren’t simply free fields, as for asymptotically free QCD Wilczek ; Gross . A well known exception is ${\cal N}=4$ SUSY Yang-Mills theory. For marginal perturbations, such as standard generalizations of $\phi^{4}$ for the Higgs potential, the renormalization group (RG) flows are rather limited in scope in exactly 4 spacetime dimensions. For scalar fields, with $\phi^{4}$ interactions, using the epsilon expansion in $D=4-\epsilon$ dimensions, one can find CFT’s in 3 spacetime dimensions such as the Wilson-Fisher fixed points, which have irrational anomalous dimensions WilsonFisher . In fact, this is at the heart of the so-called hierarchy problem, namely that straightforward RG arguments would indicate a much higher mass for the Higgs boson since there is no known ultra-violet (UV) fixed point. [^2] Without a known UV completion of the Higgs sector, there may even be higher dimension operators that make the theory non-unitary. This article will eventually scrutinize the idea that all quantum field theories (QFT’s) begin and end at RG fixed point CFT’s, at least for non-unitary theories.
@@
By contrast, in 2D one has a large toolkit of algebraic structures to study CFT’s, in particular the Virasoro algebra and affine Lie algebra symmetry of WZW models BPZ ; WittenWZW ; KZ . For the purposes of this article, we should point out that in the classification of minimal models of CFT based on the Virasoro algebra, one finds many more non-unitary theories than unitary, all with real eigenvalues of the hamiltonian, and many of them have important physical applications, such as the Lee-Yang edge singularity. Here the non-unitarity is manifested in negative norm states, rather than non-hermiticity of the hamiltonian. We would also like to stress that constructions of these CFT’s often requires projecting onto unitary sub-Hilbert spaces of non-unitary theories FQS . For instance, in the Coulomb gas method one must project the Hilbert space of a free boson to the proper space of the minimal model CFT DotsenkoFateev , and this can be described as a BRST procedure BernardFelder , or a reduction based on the SU(2) quantum group. In this way one can describe both the unitary and non-unitary minimal models of CFT in 2D. This procedure is similar to the BRST treatment of gauge theories which consistently projects out ghosts. We should also mention that the unitary CFT’s can be understood as coset reductions of current-algebras GKO , so that current algebras play a fundamental role in the classification of CFT’s.
It would certainly be nice to have such a 2D playground in 4D, and we hope this article opens such a door, however narrow it may turn out to be, based on the model we recently defined ALDolls . There we constructed some new kinds of marginal perturbations of scalar fields in 4D that have a rich structure of RG flows, at least to 1-loop. The original goal of that work was to define Higgs-like models with more interesting RG flows than the usual ones. A line of ultraviolet (UV) or infrared (IR) fixed points, and perhaps more interestingly, some cyclic RG flows were found, but based only on the 1-loop approximation. [^3] The possibility of cyclic RG flows was considered long ago in Wilson’s pioneering work on the RG KWilson . The nature of these marginal operators are such that they satisfy an interesting operator product expansion with an SU(2) Lie-algebraic structure, similar to current algebras in 2D KZ . These models can be defined for an arbitrary Lie algebra, however with applications to Higgs physics in mind, we specialized to ${\rm SU(2)}$. This algebraic structure leads to simple formulas for the 1-loop beta functions. Since the cyclic RG flows require flowing to infinite values of the couplings, it is necessary to check if the 1-loop behavior persists to higher orders. This is main subject of this paper, which can be viewed as Part II of the RG aspects of our previous work ALDolls . We will show that the 1-loop features do indeed persist non-perturbatively in a very interesting, non-trivial manner.
The regime of couplings where SU(2) is broken to U(1) is our main interest, which has 2 independent couplings. It turns out that one direction in the 2 coupling constants of our model corresponds to a line of fixed points, which are new CFT’s in 4D which we develop further below. In ALDolls we studied spontaneous symmetry breaking of SU(2) to U(1) and speculated that the cyclic regime could potentially resolve the hierarchy problem for Higgs physics. Furthermore, we found that the model has an infinite number of vacuum expectation values with Russian Doll scaling. We speculated that this could potentially explain the origin of families in the Standard Model, wherein each family is one of an infinite number of Russian Dolls which appear as resonances in scattering. The interplay between the Higgs mechanism and cyclic RG flows is clearly something worth developing further than in ALDolls , however in this article we leave this more speculative issue aside and focus solely on providing firm foundations for the quantum field theories involved based on their algebraic structures, and study their fixed points and RG flows, which are interesting in and of themselves.
As defined in ALDolls and reviewed below, the hamiltonian $H$ for our models is not hermitian, but pseudo-hermitian, which makes the model non-unitary. This is not an issue for applications to statistical mechanics in 4 spatial dimensions. For quantum mechanics in $3+1$ spacetime dimensions one needs additional positivity constraints related to unitarity. This doesn’t necessarily entail an unruly hornets’ nest if the hamiltonian has some additional algebraic structure. On the other hand, if one has merely the statement $H^{\dagger}\neq H$, then there is not even a nest to speak of. Let us mention that there has recently been renewed and growing interest in non-unitary minimal models in 2D and RG flows between them MussardoLG ; Tanaka ; KlebanovMM ; Negro . In BL38 hamiltonians satisfying $H^{\dagger}={\cal K}H{\cal K}^{\dagger}$, with ${\cal K}$ unitary, were classified according to ${\cal C},{\cal P},{\cal T}$ discrete symmetries, and 38 universality classes of random hamiltonians were obtained, extending the 3-fold classifications of Wigner and Dyson based on time-reversal symmetry, and the 10-fold classifications of Altland-Zirnbauer which included particle-hole symmetry Dyson ; AltlandZirnbauer . These classes have seen many applications to open quantum systems. The models considered in this paper are a special case where
@@
We will argue below that this structure leads to a consistent unitary quantum mechanics if one uses the operator ${\cal K}$ to project onto positive norm states. In fact pseudo-hermitian hamiltonians were proposed as consistent extensions of quantum mechanics long ago by Pauli Pauli . More recently pseudo-hermitian quantum mechanics has been developed in detail by Mostafazadeh and others Mosta1 ; Mosta2 ; Mosta3 , and is by now a well-established and highly cited framework, although one still expects some resistance. Pseudo-hermiticity as defined by (1) is not the same as ${\cal P}{\cal T}$ symmetric quantum mechanics, which also has real eigenvalues if the ${\cal P}{\cal T}$ symmetry is unbroken BenderReview .
@@
Let us summarize the main results and organization of this article. In the next section we review the definition of the models in ALDolls . In Section III we address the non-unitarity issue and propose how to project onto positive norm states which leads to a consistent quantum mechanics, with a unitary time evolution of positive probabilities. In Section IV we compute the RG beta functions to 3-loops using just the OPE structure for fully anisotropic SU(2) perturbations, which has 3 independent couplings. In Section V we specialize to SU(2) broken to U(1), which involves only 2 couplings. Based on the structure of the 3-loop results we propose beta functions to all orders. In Section VI we study the RG flows in the resulting 2-coupling parameter space to all orders in the couplings. This is made possible by a certain RG invariant $Q$ where RG flow trajectories are constant $Q$ contours, and also a strong-weak coupling duality $g\to 1/g$ of the beta functions. The resulting picture of RG flows presents itself as a kind of master flowchart for 2 couplings, since it includes new fixed point CFT’s, marginally relevant or irrelevant, asymptotic freedom, massless flows between non-trivial fixed points, and alas cyclic RG flows. [^4] This master flowchart is natural if one starts with a single coupling, which has a marginally relevant and irrelevant direction. Opening up this flow with an additional coupling that preserves this flow along the diagonals leads to this kind of flowchart. We refer to the figures below. For the cyclic flows, the RG period $\lambda$ is a simple function of the RG invariant $Q$, namely $\lambda=2\pi/\sqrt{Q}$. The existence of a regime of couplings where the RG is cyclic calls into question the commonly accepted paradigm that every QFT begins or ends at a fixed point, and we comment on this in that section. By taking one of the couplings to be imaginary, we also find massless flows between different fixed points along the critical line. The latter are 4D versions of flows between non-unitary minimal models in 2D which have received much attention recently MussardoLG ; Tanaka ; KlebanovMM ; Negro .
@@ II Definition of the models
Our models are defined as certain marginal operator perturbations of a conformal field theory consisting of two free Higgs doublets. Introduce two independent ${\rm SU(2)}$ doublets of complex bosonic spin-0 fields, $\Phi_{i}$ and $\widetilde{\Phi}_{i}$, ${i=1,2}$ and their usual hermitian conjugates $\Phi^{\dagger}$, $\widetilde{\Phi}^{\dagger}$. Under ${\rm SU(2)}$ transformations, $\Phi(x)\to U\Phi(x)$, where $U$ is a $2\times 2$ unitary ${\rm SU(2)}$ group matrix acting on the “i” indices in the above equation and the same for $\widetilde{\Phi}$. The symmetry is actually $U(2)$ which allows an additional U(1) “hypercharge” however this will not be essential to this article. The free action is the standard one for free massless bosonic fields.
${\cal S}_{0}=\int d^{4}x\left(\partial_{\mu}\Phi^{\dagger}\partial^{\mu}\Phi+\partial_{\mu}\widetilde{\Phi}^{\dagger}\partial^{\mu}\widetilde{\Phi}\right),$ (2)
@@
$[a_{\pm}({\bf{k}}),a^{\dagger}_{\pm}({\bf{k}}^{\prime})]=\delta^{(3)}({\bf{k}}-{\bf{k}}^{\prime}),~{}~{}~{}~{}[a_{\pm}({\bf{k}}),a^{\dagger}_{\mp}({\bf{k}}^{\prime})]=0,$ (4)
@@
Comparing with (11), we can interpret the operators $a_{\pm}^{\dagger}$ as creating states that are even or odd under ${\cal P}{\cal T}$. It’s important to point out that unlike ${\cal P},{\cal T}$, the operator ${\cal K}$ does not act locally on fields.
@@
$J^{a}(x)J^{b}(y)=-\dfrac{2\kappa\delta^{ab}}{16\pi^{4}|x-y|^{4}}-\dfrac{if^{abc}}{4\pi^{2}|x-y|^{2}}\,J^{c}(y)+\ldots~{}~{}~{}(\kappa=1),$ (18)
@@
$[\sigma^{a},\sigma^{b}]=if^{abc}\sigma^{c},~{}~{}~{}~{}~{}f^{abc}=2\epsilon^{abc},$ (19)
with $\epsilon^{abc}$ the completely anti-symmetric tensor with $\epsilon^{123}=1$. In the above equation $\kappa=1$, however we introduced $\kappa\neq 1$ for reasons that will become clear below. For ${\rm SU(2)}$ there is no sum over “$c$” in the equation (18) and this will simplify some calculations below. The extra minus sign for $\Phi^{\dagger_{\scriptscriptstyle\kappa}}$ in (12) is what leads to the commutator in (19) and the resulting structure constant $f^{abc}$ in the OPE (18). As we will see, the beta functions can be determined from the above OPE.
@@
${\cal O}^{a}(x)=J^{a}(x)\widetilde{J}^{a}(x),$ (20)
(no sum on $a$). The models are then defined by the action

### v3

# Non-perturbative renormalization group for pseudo-hermitian scalar fields in 4D

@@ Abstract
We define a model of 2 coupled SU(2) doublets of scalar fields in $4$ spacetime dimensions which have a rich structure of renormalization group (RG) flows to 1-loop when the SU(2) is broken to U(1). The model is pseudo-hermitian, $H^{\dagger}={\cal K}H{\cal K}^{\dagger}$ with ${\cal K}^{\dagger}{\cal K}={\cal K}^{2}=1$, which makes it non-unitary, however in a very specific manner with some desirable properties. We compute the beta functions to 3 loops from the operator product expansion and show that the 1-loop structure of flows persists to higher orders. For SU(2) broken to U(1), we conjecture a beta function to all orders. The flows can be extended to large coupling using a strong-weak coupling duality $g\to 1/g$. One finds a line of fixed points which are non-unitary conformal field theories in 4 spacetime dimensions that were previously unknown. We also find massless flows between 2 non-trivial fixed points, and a regime with a cyclic RG flow, which is allowed since the model is non-unitary. For the flows between fixed points on the critical line, we compute the anomalous dimensions of the perturbations in the UV and IR, and identify some special points where anomalous dimensions are rational numbers.
@@ I Introduction
There is a conspicuous shortage of well-understood conformal field theories (CFT’s) in 4 spacetime dimensions that aren’t simply free fields, as for asymptotically free non-abelian gauge theory such as QCD Wilczek ; Gross . A well known exception is ${\cal N}=4$ SUSY Yang-Mills theory. For marginal perturbations, such as standard generalizations of $\phi^{4}$ for the Higgs potential, the renormalization group (RG) flows are rather limited in scope in exactly 4 spacetime dimensions. In fact, this is at the heart of the so-called hierarchy problem, namely that straightforward RG arguments would indicate a much higher mass for the Higgs boson since there is no known ultra-violet (UV) fixed point. [^2] Without a known UV completion of the Higgs sector, there may even be higher dimension operators that make the theory non-unitary. For scalar fields, with $\phi^{4}$ interactions, using the epsilon expansion in $D=4-\epsilon$ dimensions, one can find CFT’s in 3 spacetime dimensions such as the Wilson-Fisher fixed points, which have irrational anomalous dimensions WilsonFisher . This article will eventually scrutinize the idea that all quantum field theories (QFT’s) begin and end at RG fixed point CFT’s, since, as we will see, our model also possesses cyclic RG flows in addition to typical flows between fixed points, which are allowed since our model is non-unitary.
@@
By contrast, in 2D one has a large toolkit of algebraic structures to study CFT’s, in particular the Virasoro algebra and affine Lie algebra symmetry of WZW models BPZ ; WittenWZW ; KZ . For the purposes of this article, we should point out that in the classification of minimal models of CFT based on the Virasoro algebra, one finds many more non-unitary theories than unitary, all with real eigenvalues of the hamiltonian, and many of them have important physical applications, such as the Lee-Yang edge singularity. Let us also mention that there has recently been renewed and growing interest in non-unitary minimal models of CFT in 2D and RG flows between them MussardoLG ; Tanaka ; KlebanovMM ; Negro ; Ravanini . Here the non-unitarity is manifested in negative norm states, rather than non-hermiticity of the hamiltonian. We would also like to stress that constructions of the unitary CFT’s often requires projecting onto unitary sub-Hilbert spaces of non-unitary theories FQS . For instance, in the Coulomb gas method one must project the Hilbert space of a free boson to the proper space of the minimal model CFT DotsenkoFateev , and this can be described as a BRST procedure BernardFelder , or a reduction based on the SU(2) quantum group. In this way one can describe both the unitary and non-unitary minimal models of CFT in 2D. This procedure is similar to the BRST treatment of gauge theories which consistently projects out the ghosts, i.e. the negative norm states. We should also mention that the unitary CFT’s can be understood as coset reductions of current-algebras GKO , so that current algebras play a fundamental role in the classification of CFT’s. It would certainly be nice to have such an algebraic formulation of CFT’s in 4D, and we hope this article opens such a door, however narrow it may turn out to be.
Let us also mention that it has been proposed that for a potential dS/CFT correspondence, the CFT is expected to be non-unitary Strominger , and symplectic fermions have been considered in this context LeClairNeubert . As we will see, there are very strong parallels between the pseudo-hermitian bosonic scalar model we study in this article and self-interacting symplectic fermions since the non-unitarity is of a very similar nature. In fact our model can be viewed as a bosonic version of the symplectic fermion. The non-unitarity of symplectic fermions has been studied in several articles Chengdu ; Ryu ; CYLee where it has been argued that the interacting model has a consistent generalized unitary time evolution at least for 2-body to 2-body scattering CYLee .
The models considered here were first defined in ALDolls where RG flows were studied to one loop. In the present article we only consider the model in 4 dimensional Euclidean space, which is sufficient for potential applications to statistical mechanics in 4 spatial dimensions. If one is interested in applications to quantum mechanics, specifically quantum field theory in $1+3$ spacetime dimensions, then one must address the non-unitarity of the model since there are additional positivity constraints. In the present work, we mainly set this issue aside since we are more interested in the RG flows and the resulting non-unitary CFT’s which are interesting in their own right. Nevertheless, let us make a few comments on the specific nature of the non-unitarity for our model in order to clarify the issue. Our hamiltonian below is pseudo-hermitian
@@
This doesn’t necessarily entail an unruly hornets’ nest if the hamiltonian has the above additional algebraic structure beyond simply the statement that $H^{\dagger}\neq H$. In BL38 hamiltonians satisfying (1), were classified according to ${\cal C},{\cal P},{\cal T}$ discrete symmetries, and 38 universality classes of random hamiltonians were obtained, extending the 3-fold classifications of Wigner and Dyson based on time-reversal symmetry, and the 10-fold classifications of Altland-Zirnbauer which included particle-hole symmetry Dyson ; AltlandZirnbauer . These classes have seen many applications to open quantum systems. In fact pseudo-hermitian hamiltonians were proposed as consistent extensions of quantum mechanics long ago by Pauli Pauli . More recently pseudo-hermitian quantum mechanics has been developed in detail by Mostafazadeh and others Mosta1 ; Mosta2 ; Mosta3 , and is by now a well-established and highly cited framework. Pseudo-hermiticity as defined by (1) is not the same as ${\cal P}{\cal T}$ symmetric quantum mechanics, which also has real eigenvalues if the ${\cal P}{\cal T}$ symmetry is unbroken BenderReview . Pseudo-hermitian hamiltonians have several desirable features that parallel the usual quantum mechanics, and we list a few. (i) Pseudo-hermitian operators $A$ satisfying $A^{\dagger}={\cal K}A{\cal K}$, such as the hamiltonian, have real eigenvalues. (ii) If one defines a modified metric on the Hilbert space which includes a $\cal K$} insertion,
$\langle\!\langle\psi^{\prime}|\psi\rangle\!\rangle\equiv\langle\psi^{\prime}|{\cal K}|\psi\rangle,$ (2)
and the $\cal K$}-conjugate
$A^{\dagger_{\scriptscriptstyle\kappa}}\equiv{\cal K}\,A^{\dagger}\,{\cal K},$ (3)
then $A^{\dagger_{\scriptscriptstyle\kappa}}$ is the proper hermitian conjugate based on the $\cal K$}-metric in (2), namely
where * denotes ordinary complex conjugation. Pseudo-hermitian operators satisfying $A^{\dagger_{\scriptscriptstyle\kappa}}=A$ all have real eigenvalues, just like the hamiltonian $H$ and thus qualify as observables. In short, the hamiltonian is hermitian with respect to the $\cal K$}-metric on the Hilbert space.  (iii) The norms of states are preserved under time evolution:
$\langle\!\langle\psi^{\prime}(t)|\psi(t)\rangle\!\rangle=\langle\!\langle\psi^{\prime}(0)|\psi(0)\rangle\!\rangle.$ (5)
For our model, the $\cal K$}-metric introduces negative norm states since ${\cal K}^{2}=1$ implies states with both ${\cal K}=\pm 1$. Although (5) implies the sum of probabilities for outcomes of measurements all add up to $1$ for all times if $|\psi(0)\rangle$ is normalized to $1$, these negative norm states imply some probabilities are negative, which obviously needs to be addressed. On the other hand, the free hamiltonian $H_{0}$ below commutes with $\cal K$}, thus negative norm states can be consistently projected out in the free theory with the projectors
$P_{\pm}=\dfrac{1\pm{\cal K}}{2},~~~~P_{+}+P_{-}=1,~~~~~P_{\pm}^{2}=P_{\pm},~~~~~P_{+}P_{-}=0,$ (6)
since ${\cal K}P_{\pm}|\psi\rangle=\pm P_{\pm}|\psi\rangle$. However since the full interacting hamiltonian $H$ below does not commute with $\cal K$}, otherwise $H^{\dagger}=H$, projecting onto positive norm states with $P_{+}$ does not lead to completely consistent quantum mechanics due to $P_{-}e^{-iHt}P_{+}|\psi\rangle\neq 0$ since $\cal K$} does not commute with $H$. Thus, starting from a positive norm state $P_{+}|\psi\rangle$, time evolution can produce negative norm states, which needs additional considerations in order to have a quantum mechanics consistent with our current interpretation of the probabilistic aspect of measurement. Based on the review of Mostafazadeh Mosta1 , given the above pseudo-hermitian properties, in principle there exists a prescription for projecting onto positive norm states with further modifications of the metric on the Hilbert space, however this procedure is model-dependent, not unique, and somewhat complicated, thus we do not pursue this here. As previously stated, we henceforth leave this issue aside and simply focus on the RG aspects of the non-unitary Euclidean QFT we define below. [^3] In the original version of our article ALDolls , the non-unitarity issue was not fully resolved, and we are preparing a revised version to address this based on generalizing what was understood recently for symplectic fermions CYLee . In that article we also speculated that the existence of families in the Standard Model of particle physics is perhaps the result of the cyclic RG we find below, wherein each family corresponds to 1 cycle of the RG with the typical Russian Doll scaling of masses, however more work is needed towards a complete model that realizes this idea.
We will define some new type of marginal perturbations of scalar fields in 4D that have a rich structure of RG flows. To 1-loop, one finds a line of ultraviolet (UV) or infrared (IR) fixed points, and perhaps more interestingly, some cyclic RG flows in the 1-loop approximation. The possibility of cyclic RG flows was considered long ago in Wilson’s pioneering work on the RG KWilson . The nature of these marginal operators are such that they satisfy an interesting operator product expansion with an SU(2) Lie-algebraic structure, similar to current algebras in 2D KZ . These models can be defined for an arbitrary Lie algebra, however we specialize to $\rm SU(2)$}. This algebraic structure leads to simple formulas for the 1-loop beta functions. Since the cyclic RG flows require flowing to infinite values of the couplings, it is necessary to check if the 1-loop behavior persists to higher orders, and this is the main subject of this article. We will show that the 1-loop features do indeed persist non-perturbatively in a very interesting, non-trivial manner.
@@
Let us summarize the main results and organization of this article. In the next section we review the definition of the models. In Section III we compute the RG beta functions to 3-loops using just the OPE structure for fully anisotropic SU(2) perturbations, which has 3 independent couplings. In Section IV we specialize to SU(2) broken to U(1), which is our main interest, and involves only 2 couplings. Based on the structure of the 3-loop results we propose beta functions to all orders. In Section V we study the RG flows in the resulting 2-coupling parameter space to all orders in the couplings. This is made possible by a certain RG invariant $Q$ where RG flow trajectories are constant $Q$ contours, and also a strong-weak coupling duality $g\to 1/g$ of the beta functions. The resulting picture of RG flows presents itself as a kind of master flowchart for 2 couplings, since it includes new fixed point CFT’s, marginally relevant or irrelevant, asymptotic freedom, massless flows between non-trivial fixed points, and alas cyclic RG flows. [^4] This master flowchart is natural if one starts with a single coupling, which has a marginally relevant and irrelevant direction. Opening up this flow with an additional coupling that preserves this flow along the diagonals leads to this kind of flowchart. We refer to the figures below. For the cyclic flows, the RG period $\lambda$ is a simple function of the RG invariant $Q$, namely $\lambda=2\pi/\sqrt{Q}$. The existence of a regime of couplings where the RG is cyclic calls into question the commonly accepted paradigm that every QFT begins or ends at a fixed point, and we comment on this in that section. By taking one of the couplings to be imaginary, we also find massless flows between different fixed points along the critical line. The latter are 4D versions of flows between non-unitary minimal models in 2D which have received much attention recently MussardoLG ; Tanaka ; KlebanovMM ; Negro .
@@ II Definition of the models
Our models are defined as certain marginal operator perturbations of a conformal field theory consisting of two scalar field doublets. Introduce two independent $\rm SU(2)$} doublets of complex bosonic spin-0 fields, $\Phi_{i}$ and $\widetilde{\Phi}_{i}$, i=1,2{i=1,2} and their usual hermitian conjugates $\Phi^{\dagger}$, $\widetilde{\Phi}^{\dagger}$. Under $\rm SU(2)$} transformations, $\Phi(x)\to U\Phi(x)$, where $U$ is a $2\times 2$ unitary $\rm SU(2)$} group matrix acting on the “i” indices in the above equation and the same for $\widetilde{\Phi}$. The symmetry is actually $U(2)$ which allows an additional U(1) “hypercharge” however this will not be essential to this article. The free action is the standard one for free massless bosonic fields.
[…]
@@
$a_{\pm}({\bf{k}}),a^{\dagger}_{\pm}({\bf{k}}^{\prime})]=\delta^{(3)}({\bf{k}}-{\bf{k}}^{\prime}),~~~~[a_{\pm}({\bf{k}}),a^{\dagger}_{\mp}({\bf{k}}^{\prime})]=0,$ (9)
@@
It’s important to point out that unlike ${\cal P},{\cal T}$, the operator $\cal K$} does not act locally on fields.
@@
$J^{a}(x)J^{b}(y)=-\dfrac{2\kappa\delta^{ab}}{16\pi^{4}|x-y|^{4}}-\dfrac{if^{abc}}{4\pi^{2}|x-y|^{2}}\,J^{c}(y)+\ldots~~~({\rm with}~\kappa=1),$ (23)
@@
$\sigma^{a},\sigma^{b}]=if^{abc}\sigma^{c},~~~~~f^{abc}=2\epsilon^{abc},$ (24)
with $\epsilon^{abc}$ the completely anti-symmetric tensor with $\epsilon^{123}=1$. In the above equation $\kappa=1$, however we introduced $\kappa\neq 1$ for book-keeping reasons that will become clear below, as the $\kappa$ order counts loops for contributions to the beta function. For $\rm SU(2)$} there is no sum over “cc” in the equation (23) and this will simplify some calculations below. The extra minus sign for $\Phi^{\dagger_{\scriptscriptstyle\kappa}}$ in (17) is what leads to the commutator in (24) and the resulting structure constant $f^{abc}$ in the OPE (23). If $J^{a}$ was instead defined as $\phi^{\dagger}\sigma^{a}\phi$, it would not satisfy the OPE in (23). As we will see, the beta functions can be determined from the above OPE (23).
@@
${\cal O}^{a}(x)=J^{a}(x)\widetilde{J}^{a}(x).$ (25)
The models are then defined by the action